Post Card Depot, an large retailer of post cards, orders 6,317,982 post cards per year from its manufacturer. Post Card Depot plans on ordering post card 19 times over the next year. Post Card Depot receives the same number of post cards each time it orders. The carrying cost is $0.14 per post card per year. The ordering cost is $337 per order.

What is the annual carrying costs of post card inventory (round the answer to two decimal places)?
Answer #1

ordering cost = 19*337 = 6403

Carrying costs = (6317982/19) /2 * 0.14= 332525/2 *0.14

= 166263 *0.14

= 23276.82

Total cost = ordering cost + carrying cost

= 23276.82 + 6403

= 29679.82